OpenStudy (anonymous):

Find the derivative of f(x) = 5x + 9 at x = 2.

OpenStudy (anonymous):

If x=2 then\[f(2)=7+9=16\]The derivative of a constant is zero so,\[f'(2)=0\]

OpenStudy (anonymous):

I meant: \[f(2)=10+9=19\]\[f'(2)=0\]

OpenStudy (phoenixfire):

Would you not find the derivative and then pass in 2 as x. \[f(x)=5x+9\] \[f \prime (x) = 5\] Therefore, \[f \prime (2) = 5\]